What should be the rate of a telephone line that can carry 16 voice channels? show work
Answer #1

No of channel = 16

Voice Frequency range = 4 kHz

Hence sampling rate applying Nyquest Theorem = 2 * 4 = 8 kHz

Each sample = 8 bit
Data rate in 1 channel = sampling rate * (no of bit in a sample) = 8*8 = 64 Kbps


Rate of telephone line: No of channel * 64 Kbps = 1024 Kbps = 1 Mbps

Ans: Hence the rate of telephone line is 1Mbps for 16 voice or audio channel.